OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = n (6n^2 - 3n - 1)/2

Nnesha (nnesha):

hey do you have an idea how to start this ??

Nnesha (nnesha):

`1st)` substitute n for 1 to check is the L.H.S = R.H.S ? if the statement is true then next step is `2nd)` assume it is true for n=k (substitute n for k) this step is called " induction assumption' `3rd)` substitute n for k+1 and we want to show the statement is true for n= k+1 based on the 2nd step assumption

Nnesha (nnesha):

just tag me when u rready

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

I substituted n with 1 on the left hand side to get \[\frac{ 1[6(1)^2-3(1)-1] }{ 2 }\]

OpenStudy (anonymous):

Which would then make the LHS = 1

OpenStudy (anonymous):

I dont know what to do with the RHS

OpenStudy (anonymous):

I mean... I did that to the RHS and dont know what to do with the LHS

Nnesha (nnesha):

for first step (3n - 2)^2 = n (6n^2 - 3n - 1)/2 you need to see if both sides are equal when u substitute n for 1

Nnesha (nnesha):

\[\large\rm (3(1)-2)^2=\frac{ 1[6(1)^2-3(1)-1] }{ 2 }\] both sides are equa l? then move to the next step

OpenStudy (anonymous):

substitute n for k???

Nnesha (nnesha):

yes right that would be the 2nd step and have to assume that n = k is true

OpenStudy (anonymous):

What is k?

Nnesha (nnesha):

k is just a variable that represent the highest term kth term

Nnesha (nnesha):

not the highest i mean one of the term

OpenStudy (anonymous):

So...\[(3k-2)^2=\frac{ 1[6(k)^2-3(k)-1]}{ 2 }\]

OpenStudy (anonymous):

\[(3k−2)2=k[6(k)2−3(k)−1]2\] *

OpenStudy (anonymous):

oops...... \[(3k−2)^2=\frac{k[6(k)^2−3(k)−1] }{ 2 }\]

Nnesha (nnesha):

\[\rm 1^2 + 4^2 + 7^2 + ... + (3\color{red}{k} - 2)^2 = \frac{\color{Red}{k} (6\color{Red}{k}^2 - 3\color{Red}{n} - 1)}{2}\] for 2nd step keep the first terms bec this is the assumption we need for 3rd step

Nnesha (nnesha):

we assumed that n =k is true now move on to the 3rd step substitute all n's with k+1

OpenStudy (anonymous):

is the 3n supposed to be 3n or 3k?

Nnesha (nnesha):

oh ye iforgot to change that

OpenStudy (anonymous):

\[1^2+4^2+7^2+...+(3(k+1)−2)^2=\frac{ (k+1)[6(k+1)^2−3(k+1)−1] }{ 2 }\] This is starting to look a little bit less pleasant

Nnesha (nnesha):

\[\rm 1^2 + 4^2 + 7^2 + ... + (3\color{red}{k} - 2)^2 = \frac{\color{Red}{k} (6\color{Red}{k}^2 - 3\color{Red}{k} - 1)}{2}\] now substitute n for k+1 and we should leave the last term which is (3k-2)^2 at left side \[\rm 1^2 + 4^2 + 7^2 + ... + \color{blue}{(3\color{blue}{k} - 2)^2}+(3[\color{red}{k+1}]-2)^2 = \frac{\color{Red}{k+1} (6\color{Red}{[k+1]}^2 - 3\color{Red}{[k+1]} - 1)}{2}\]

Nnesha (nnesha):

hahah ik but it will be fun!!

Nnesha (nnesha):

that's correct but we need to leave (3k-2)^2 at left side and add (3[k+1]-2)^2 because here we gotta use the assumption

Nnesha (nnesha):

\[\rm \color{orange }{1^2 + 4^2 + 7^2 + ... + (3\color{orange}{k} - 2)^2} = \frac{\color{Red}{k} (6\color{Red}{k}^2 - 3\color{Red}{k} - 1)}{2}\] according to the 2nd assumption (the orange side ) is equal to k(6k^2+3k-1)/2 so can substitute (the orange part) with k(6k^2+3k-1)/2 \[\rm \color{orange}{\frac{k(6k^2+3k-1)}{2}}+(3[\color{red}{k+1}]-2)^2 = \frac{\color{Red}{k+1} (6\color{Red}{[k+1]}^2 - 3\color{Red}{[k+1]} - 1)}{2}\]

Nnesha (nnesha):

make sense ? if no feel free to ask question :)

OpenStudy (anonymous):

Attempting to solve. Dont think I know how.

Nnesha (nnesha):

well did you understand the last part ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I understood the substitution part

Nnesha (nnesha):

alright

Nnesha (nnesha):

remember the PEMDAS rule parentheses first then exponent try to solve :=))

OpenStudy (anonymous):

LHS= \[\frac{ 6k^3+3k^2-k }{ 2 }+(3[k+1]-2)^2\]

Nnesha (nnesha):

good so far :=))

OpenStudy (anonymous):

\[(3[k+1]−2)^2\] becomes \[9k^2-1\]?

OpenStudy (anonymous):

\[\frac{ 6k^3+3k^2−k }{ 2 } + 9k^2 -1\]

OpenStudy (anonymous):

Not sure about the other side

OpenStudy (anonymous):

I'll try to finish this tomorrow...

Nnesha (nnesha):

ugh sorry i was afk let me chec k

Nnesha (nnesha):

hmmm well it can't be negative one it's not difference of squares \[a^2-b^2=(a-b)(a+b)\]

Nnesha (nnesha):

\[(a+b)^2=a^2+2ab+b^2\] first distribute bracket by 3

Nnesha (nnesha):

alright i'll be free tomorrw :=))

OpenStudy (anonymous):

@Nnesha Do you know if it is true or false?

OpenStudy (anonymous):

The world may never know...